# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ B.1 Global minimum variance portfolio
Figure 10(a)-10(c) present the results obtained using historical data of varying lengths with $q\in\mathcal{Q}_{1}$, and Figure 10(d)-10(f) use $q\in\mathcal{Q}_{2}$. From Figure 10, we can make a few conclusions. Firstly, with no regularization ($q=0$, the blue curve), the out-of-sample volatility is very high. So adding proper regularization to protect the condition of the sample covariance is absolutely necessary. Secondly, as shown in Figure 10, for each of the 6 cases, if we fix one choice $q\in\mathcal{Q}$ for all testing months, for most months $q=\min_{q\in\mathcal{Q}}q$ (the orange curve) achieves the smallest out-of-sample volatility, but occasionally $q=\max_{q\in\mathcal{Q}}q$ (the green curve) outperforms $q=\min_{q\in\mathcal{Q}}q$. If the volatility happens to be monotone in the range of $q$ values in $\mathcal{Q}$, the boxplot will be in the middle between the orange and green curves, but if the best $q$ is neither the minimum $q$ nor the maximum $q$, we also see cases that the boxplot is beneath both the orange and green curves. This indicates that applying a constant level of regularization for all months such as $q=\min_{q\in\mathcal{Q}}q$ to the portfolio may not be suitable for all time periods. So if we want to update $q$ values from month to month, it is important to have a in-sample method to estimate the out-of-sample volatility. Thirdly, from all 6 cases in Figure 10, we see that the portfolios with optimized $q^{*}$ (the red triangles), automatically updated every month, consistently exhibit a smaller standard deviation regardless of the candidate sets and the varying lengths of historical data. So we can conclude that our method can consistently and flexibly select an appropriate $q\in\mathcal{Q}$ based on the data itself, leading to smaller out-of-sample variances of the global minimum variance portfolio and demonstrating significant advantages over naively choosing a fixed $q\in\mathcal{Q}$ for all time points. The last side observation is that the volatility level jumps to a higher level starting from the label 201707, which actually represents July 2017 to June 2020, the first time period covering the COVID outbreak.
@@ B.2 Calibrated Models with Real Data
In the previous experiments for mean-variance portfolios, we made the assumption that the mean vector ${\bm{\mu}}$ in S&P500 is known. However, in reality, it is typically unknown. In this section, we hope to look into how our methods are affected when ${\bm{\mu}}$ is unobserved. To this end, we use real data to calibrate a model for ${\bm{\mu}}$ and then use the calibrated model to simulate synthetic data. Here are the steps of this experiment.
@@
    $\begin{split}&\check{{\bm{\mu}}}_{j}=\rho\cdot{\bm{\mu}}_{j}+\sqrt{1-\rho^{2}}\cdot\mathcal{N}(0,\sigma^{2}_{{\bm{\mu}}_{j}}\mathbf{I}),\\ &\check{{\bm{\Sigma}}}_{j}=0.95\cdot\widehat{\bm{\Sigma}}_{j}+\sqrt{1-0.95^{2}}\cdot\widetilde{\bm{\Sigma}}_{j}.\end{split}$ (B.2)
@@
    To use the calibrated model, we assume that the mean vector ${\bm{\mu}}_{j}$ is known and used in portfolio construction, but the actual true mean of the returns for the month $j$ is $\check{{\bm{\mu}}}_{j}$ which is unknown. We use different values of $\rho$ to generate different calibrated models, where a larger $\rho$ reflects a higher signal level of predicting $\check{{\bm{\mu}}}_{j}$ using ${\bm{\mu}}_{j}$. The allocation vector for the testing month $j$ is then constructed as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}_{j}$, and the remaining steps follow the same procedure in Section 5.1.
@@
We present experimental results for $\rho=0.05$, $0.2$, and $0.4$ in three rows of plots respectively in Figure 11. As indicated in equation (B.2), smaller values of $\rho$ result in larger perturbations on the mean vector ${\bm{\mu}}_{j}$. $\rho=0.05$, $0.2$, and $0.4$, the mean cosine values of the angle between ${\bm{\mu}}_{j}$ and $\check{{\bm{\mu}}}_{j}$ are $0.0428$, $0.1368$, and $0.2605$, respectively, with corresponding standard deviations of $0.2676$, $0.2649$, and $0.2562$. We observed that as $\rho$ increases, the Sharpe ratio for the portfolio gradually increases. This makes intuitive sense as higher prediction power leads to better portfolio out-of-sample performance. Specifically, for $\rho=0.2$ and $0.4$, the optimized $q^{*}$ based on our in-sample SR estimator always displays the best performance among the candidate $q$ values in $\mathcal{Q}=\mathcal{Q}_{1}$ or $\mathcal{Q}=\mathcal{Q}_{2}$. For $\rho=0.05$, we may not always achieve the best out-of-sample SR due to the low signal level ($\check{{\bm{\mu}}}_{j}$ becomes nearly orthogonal to ${\bm{\mu}}_{j}$). However, we still see comparable performances.
@@ Appendix C Supporting Lemmas
In this section, we will present several lemmas that play a crucial role in determining the asymptotic behavior of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$. These lemmas are instrumental in our analysis and provide valuable insights into the problem at hand.
@@
Given $\mathbf{X}\in\mathbb{R}^{n\times p}$ with $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}$ where the elements of $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are i.i.d with zero mean, variance 1 and finite $8+\varepsilon$ order moment for some $\varepsilon>0$, ${\bm{\Sigma}}^{\frac{1}{2}}\in\mathbb{R}^{p\times p}\geq 0$ is a non negative Hermitian matrix with $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ bounded. For any deterministic $\mathbf{A}\in\mathbb{R}^{p\times p}$ with bounded trace norm (${\mathrm{tr}}(\sqrt{\mathbf{A}\mathbf{A}^{\top}})\leq C$ with respect to $n$) and Hermitian positive matrix $\mathbf{S}\in\mathbb{R}^{p\times p}$, define
@@
and assume that $p,n$ tends to infinity with proportion $p/n\to c>0$. Then for $z\in\mathbb{C}-(-\infty,0]$,
@@
Lemma C.6 established the key asymptotic behavior in Case 2 in Assumption 2.6. However, a notable distinction in Case 3 in Assumption 2.6 is the presence of $p/n\to c\geq 1$. Consequently, the proof technique utilized in Lemma C.6 is no longer applicable due to the inability to uniformly bound $|f^{\prime}_{n}(z)|$ by a trivial constant. Nevertheless, we can still address the case $z=0$ if we make the additional assumption that the number of eigenvalues $K$ tending to infinity remains fixed and $\lambda_{1}/\lambda_{K}\leq C_{\lambda}$:
@@
$|f_{n}^{\prime}(z)|={\mathrm{tr}}\bigg{[}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2}\bigg{]}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\leq\frac{\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathrm{tr}}}}{\lambda_{\min}(\mathbf{Z}^{\top}\mathbf{Z}/n)}\leq\frac{4(\|{\bm{\Sigma}}^{-1}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{A}\|_{{\mathrm{tr}}})}{(1-\sqrt{c})^{4}}.$
The second inequality holds almost surely for large enough $n$, by the Bai-Yin theorem (Bai and Yin, 1993). As its derivatives are bounded, the sequence $f_{n},n=1,2,3,\ldots$ is equicontinuous, and by the Arzela-Ascoli theorem, we deduce that $f_{n}$ converges uniformly to its limit. By the Moore-Osgood theorem, we can exchange limits (as $n,p\rightarrow\infty$ and $z\rightarrow 0^{+}$).
@@
$f_{n}^{\prime}(z)$ $={\mathrm{tr}}\bigg{[}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{]}\mathbf{A},$
if we can prove that for any fixed vector $\mathbf{u}$ with $\|\mathbf{u}\|=1$,
@@
almost surely for large enough $n$. Here, $\lambda_{\mathbf{A},i}$ is the eigenvalue of matrix $\mathbf{A}$ and ${\bm{\xi}}_{\mathbf{A},i}$ is the corresponding eigenvector. Hence, the main target now is to deduce the uniform bound of (C.6). We separate the proof of the uniform bound of (C.6) into the case $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$ and $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$
When $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, it can be seen that
(C.6)$\leq\frac{\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{u}\|_{2}^{2}}{z^{2}}<+\infty,$
which completes the proof that (C.6) is uniformly bounded when $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$.
When $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, by Assumption 2.6, the number of eigenvalues tending to infinity remains fixed, we can separate ${\bm{\Sigma}}$ by
@@
Therefore, when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$, we have almost surely
$\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{K}\leq(z\cdot K\lambda_{j}\cdot\sqrt{3c}/\sqrt{\lambda_{j}})^{2}/\lambda_{K}\leq z^{2}\cdot K^{2}C_{\lambda}\cdot 3c\leq\frac{\lambda_{\min}(\mathbf{Q})}{6}$
for $n$ large enough. Hence when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,
@@
almost surely for all $n$ large enough. Here, the first inequality utilizes ${\widehat{\lambda}}_{K}\geq 2\lambda_{K}/3$ almost surely for large enough $n$. We prove that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,
$\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\cdot\lambda_{K}=\lambda_{K}/\lambda_{\min}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})\leq 2$ (C.11)
almost surely for all $n$ large enough. Note that $\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}^{2}/\lambda_{K}\leq\lambda_{\min}(\mathbf{Q})/6$ as long as $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$ for $n$ large enough, we conclude that
$\|{\bm{\Delta}}_{bs}\|_{{\mathop{\mathrm{op}}}}\cdot\sqrt{\lambda_{K}}\leq\lambda_{K}\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}/(\lambda_{\min}(\mathbf{Q})\cdot\sqrt{\lambda_{K}})\leq\sqrt{\frac{2}{3\lambda_{\min}(\mathbf{Q})}}$ (C.12)
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$. As for $\|{\bm{\Delta}}_{bs}\|_{{\mathop{\mathrm{op}}}}$, we have
@@
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$. Combined (C.9) with (C.10)-(C.13), we conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18cC_{\lambda}K^{2}}}$,

### v3

@@ B.1 Global minimum variance portfolio
Figure 11(a)-11(c) present the results obtained using historical data of varying lengths with $\mathbf{Q}\in\mathcal{Q}_{1}$, and Figure 11(d)-11(f) use $\mathbf{Q}\in\mathcal{Q}_{2}$. From Figure 11, we can make a few conclusions. Firstly, with no regularization ( $\mathbf{Q}={\bm{0}}$, the blue curve), the out-of-sample volatility is very high. So adding proper regularization to protect the condition of the sample covariance is absolutely necessary. Secondly, as shown in Figure 11, for each of the 6 cases, if we fix one choice $\mathbf{Q}\in\mathcal{Q}$ for all testing months, for most months $q=0.1$ (the orange curve, minimum tuning parameter) achieves the smallest out-of-sample volatility, but occasionally $q=3$ (the green curve, minimum tuning parameter) outperforms $q=0.1$. If the volatility happens to be monotone in the range of $q$ values, the boxplot will be in the middle between the orange and green curves, but if the best $q$ is neither the minimum $q$ nor the maximum $q$, we also see cases that the boxplot is beneath both the orange and green curves. This indicates that applying a constant level of regularization for all months such as $q=0.1$ (minimum tuning parameter) to the portfolio may not be suitable for all time periods. So if we want to update $\mathbf{Q}$ from month to month, it is important to have a in-sample method to estimate the out-of-sample volatility. Thirdly, from all 6 cases in Figure 11, we see that the portfolios with optimized $\mathbf{Q}^{*}$ (the red triangles), automatically updated every month, consistently exhibit a smaller standard deviation regardless of the candidate sets and the varying lengths of historical data. So we can conclude that our method can consistently and flexibly select an appropriate $\mathbf{Q}\in\mathcal{Q}$ based on the data itself, leading to smaller out-of-sample variances of the global minimum variance portfolio and demonstrating significant advantages over naively choosing a fixed $\mathbf{Q}\in\mathcal{Q}$ for all time points. The last side observation is that the volatility level jumps to a higher level starting from the label 201707, which actually represents Jul 2017 to Jun 2020, the first time period covering the COVID outbreak.
@@ B.2 Calibrated models with real data
In the previous experiments for mean-variance portfolios, we made the assumption that the mean vector ${\bm{\mu}}$ in S&P500 is known. In this section, we hope to look into how our methods are affected when ${\bm{\mu}}$ is unobserved. To this end, we use real data to calibrate a model for ${\bm{\mu}}$ and then use the calibrated model to simulate synthetic data. Here are the steps.
@@
    $\begin{split}\check{{\bm{\mu}}}_{j}=\rho\cdot{\bm{\mu}}_{j}+\sqrt{1-\rho^{2}}\cdot\mathcal{N}(0,\sigma^{2}_{{\bm{\mu}}_{j}}\mathbf{I}),\quad\check{{\bm{\Sigma}}}_{j}=0.95\cdot\widehat{\bm{\Sigma}}_{j}+\sqrt{1-0.95^{2}}\cdot\widetilde{\bm{\Sigma}}_{j}.\end{split}$ (B.2)
@@
    To use the calibrated model, we assume that the mean vector ${\bm{\mu}}_{j}$ is known and used in portfolio construction, but the actual true mean of the returns for the month $j$ is $\check{{\bm{\mu}}}_{j}$ which is unknown. We use different values of $\rho$ to generate different calibrated models, where a larger $\rho$ reflects a higher signal level of predicting $\check{{\bm{\mu}}}_{j}$ using ${\bm{\mu}}_{j}$. The allocation vector for the testing month $j$ is then constructed as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}_{j}$, and the remaining steps follow the same procedure in Section 5.1.
@@
We present experimental results for $\rho=0.05$, $0.2$, and $0.4$ in three rows of plots respectively in Figure 12. As indicated in equation (B.2), smaller values of $\rho$ result in larger perturbations on the mean vector ${\bm{\mu}}_{j}$. $\rho=0.05$, $0.2$, and $0.4$, the mean cosine values of the angle between ${\bm{\mu}}_{j}$ and $\check{{\bm{\mu}}}_{j}$ are $0.0428$, $0.1368$, and $0.2605$, respectively, with corresponding standard deviations of $0.2676$, $0.2649$, and $0.2562$. We observed that as $\rho$ increases, the Sharpe ratio for the portfolio gradually increases. This makes intuitive sense as higher prediction power leads to better portfolio out-of-sample performance. Specifically, for $\rho=0.2$ and $0.4$, the optimized $\mathbf{Q}^{*}$ based on our in-sample SR estimator always displays the best performance among the candidate $\mathbf{Q}$ in $\mathcal{Q}=\mathcal{Q}_{1}$ or $\mathcal{Q}=\mathcal{Q}_{2}$. For $\rho=0.05$, we may not always achieve the best out-of-sample SR due to the low signal level ($\check{{\bm{\mu}}}_{j}$ becomes nearly orthogonal to ${\bm{\mu}}_{j}$). However, we still see comparable performances.
B.3 Mean variance portfolio with unknown mean vector
In this section, we investigate the MV portfolio construction where ${\bm{\mu}}$ is unknown. When ${\bm{\mu}}$ is unknown, we define the MV portfolio allocation vector as $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}$. The portfolio is scaled under the assumption $\sum_{i}|\mathbf{w}_{i}|=1$, which ensures that the total exposure of the portfolio sums to one unit. Thus, the weight vector $\mathbf{w}$ can be expressed as:
$\mathbf{w}=(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}/\|(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}\|_{1}.$ (B.3)
In this experiment, we change the set $\mathcal{Q}_{1}=\{q\cdot{\rm diag}(\widehat{\bm{\Sigma}}_{pre}-\lambda_{1}{\bm{\xi}}_{1}{\bm{\xi}}_{1}^{\top}),q\in[1:30]/10\}$, where $\widehat{\bm{\Sigma}}_{pre}$ represents the sample covariance matrix of returns calculated from Jan 2004 to Dec 2008, and $\lambda_{1}$ and ${\bm{\xi}}_{1}$ are the largest eigenvalues and corresponding eigenvectors of $\widehat{\bm{\Sigma}}_{pre}$, and we still keep the set $\mathcal{Q}_{2}$. According to Theorem G.2, for each $\mathbf{Q}\in\mathcal{Q}$, we can leverage $\widehat{SR}$ to identify the optimal regularization matrix $\mathbf{Q}$ within $\mathcal{Q}_{1}$ or $\mathcal{Q}_{2}$ that maximizes the Sharpe ratio ($SR$). We present the steps for our MV portfolio analysis.
  1. 1.
    For each testing month, we observe the excess daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ is the total number of trading days with one-, two-, and four-year historical data length and $p=365$ is the number of selected stocks. Calculate the sample mean $\bm{\widehat{\mu}}$, the sample covariance matrix $\widehat{\bm{\Sigma}}$ and the portfolio weight $\mathbf{w}$ given by (B.3).
  2. 2.
    For each testing month, we run experiments for all candidate values and also consider no regularization, i.e. $\mathbf{Q}={\bm{0}}$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}\bm{\widehat{\mu}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ using the estimation in Theorem G.2. Here,
    $\mathbf{Q}^{*}=\operatorname{argmax}_{\mathbf{Q}^{*}\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bigg{)}\cdot\frac{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}-\frac{{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}}{n-{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}}}{\sqrt{\bm{\widehat{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bm{\widehat{\mu}}}}.$
  3. 3.
    We roll the procedure above for all testing months. Note that the optimal $\mathbf{Q}^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $\mathbf{Q}\in\mathcal{Q}$, $\mathbf{Q}={\bm{0}}$ or $\mathbf{Q}^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
  4. 4.
    We report the realized Sharpe ratio of daily portfolio returns over the future three years.
Figure 13: SR of mean-variance portfolios. The x-axis labels the rolling period, while the y-axis represents the out-of-sample SR of the portfolio returns every three years in the future. The blue solid line, orange dash-dot line and green dash line correspond to the SR with $q=0$, $q=0.1$ (minimum $q$) and $q=3$ (maximum $q$) respectively. The boxplot displays the Sharpe ratio for all $\mathbf{Q}\in\mathcal{Q}$, and the red triangle indicates the SR under our optimized $\mathbf{Q}^{*}\in\mathcal{Q}$.
Here we apply the sample mean and covariance matrix to construct the portfolio weight $\mathbf{w}$. Different from Section 5.1, we do not access to the mean vector ${\bm{\mu}}$ in the testing month. Figures 13(a) to 13(c) present the results for $\mathbf{Q}\in\mathcal{Q}_{1}$, while Figures 13(d) to 13(f) display the outcomes for $\mathbf{Q}\in\mathcal{Q}_{2}$. The blue curve represents the scenario where $\mathbf{Q}={\bm{0}}$. The results clearly highlight the benefits of an actively optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ for each testing month compared to a static, fixed $\mathbf{Q}\in\mathcal{Q}$ used uniformly across all months. Moreover, using the sample mean $\bm{\widehat{\mu}}$ leads to a significant drop in the Sharpe ratio compared to the scenario where ${\bm{\mu}}$ is known. This observation underscores the importance of accurate ${\bm{\mu}}$ estimation in practical settings. It suggests that relying the sample mean to estimate ${\bm{\mu}}$ may be insufficient.
@@ Appendix C Supporting Lemmas
In this section, we will present several lemmas that play a crucial role in determining the asymptotic behavior of $T_{n,1}(\mathbf{Q})$ and $T_{n,2}(\mathbf{Q})$. These lemmas are instrumental in our analysis and provide valuable insights into the problem at hand. By Assumption 2.5, readers may keep in mind that $\lambda_{\min}(\mathbf{Q})$ is bounded away from 00, and we let $c_{\mathbf{Q}}$ be a constant that is smaller than $\lambda_{\min}(\mathbf{Q})$. The special case $\mathbf{Q}={\bm{0}}$ when $c<1$ is excluded in this section, the proof of Theorem 2.7 when $\mathbf{Q}={\bm{0}}$ is simple and we directly show it in Section D.
@@
Given $\mathbf{X}\in\mathbb{R}^{n\times p}$ with $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}$ where the elements of $\mathbf{Z}\in\mathbb{R}^{n\times p}$ are i.i.d with zero mean, variance 1 and finite $8+\varepsilon$ order moment for some $\varepsilon>0$, ${\bm{\Sigma}}^{\frac{1}{2}}\in\mathbb{R}^{p\times p}\geq 0$ is a non negative Hermitian matrix with $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ bounded. For any deterministic $\mathbf{A}\in\mathbb{R}^{p\times p}$ with bounded trace norm (${\mathrm{tr}}(\sqrt{\mathbf{A}\mathbf{A}^{\top}})\leq C$ with respect to $n$) and Hermitian non-negative matrix $\mathbf{S}\in\mathbb{R}^{p\times p}$, define
@@
and assume that $p,n$ tends to infinity with proportion $p/n\to c>0$. Then for fixed $z\in\mathbb{C}-(-\infty,0]$,
@@
Lemma C.6 established the key asymptotic behavior in Case 2 in Assumption 2.6. However, a notable distinction in Case 3 in Assumption 2.6 is the presence of $p/n\to c\geq 1$. Consequently, the proof technique utilized in Lemma C.6 is no longer applicable due to the inability to uniformly bound $|f^{\prime}_{n}(z)|$ by a trivial constant. Nevertheless, we can still address the case $z=0$ if we make the additional assumption that the number of eigenvalues $K$ tending to infinity remains fixed and $\lambda_{1}/\lambda_{K}^{2}\leq C$:
@@
$|f_{n}^{\prime}(z)|={\mathrm{tr}}\bigg{[}\bigg{(}\frac{\mathbf{Z}^{\top}\mathbf{Z}}{n}+{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{Q}{\bm{\Sigma}}^{-\frac{1}{2}}+z\mathbf{I}\bigg{)}^{-2}\bigg{]}{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\leq\frac{\|{\bm{\Sigma}}^{-\frac{1}{2}}\mathbf{A}{\bm{\Sigma}}^{-\frac{1}{2}}\|_{{\mathrm{tr}}}}{\lambda_{\min}^{2}(\mathbf{Z}^{\top}\mathbf{Z}/n)}\leq\frac{4(\|{\bm{\Sigma}}^{-1}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{A}\|_{{\mathrm{tr}}})}{(1-\sqrt{c})^{4}}.$
The second inequality holds almost surely for large enough $n$, by the Bai-Yin theorem (Bai & Yin 1993) that $\lambda_{\min}(\mathbf{Z}^{\top}\mathbf{Z}/n)\geq(1-\sqrt{c})^{2}/2$ almost surely for all $n$ large enough. As its derivatives are bounded, the sequence $f_{n},n=1,2,3,\ldots$ is equicontinuous, and by the Arzela-Ascoli theorem, we deduce that $f_{n}$ converges uniformly to its limit. By the Moore-Osgood theorem, we can exchange limits (as $n,p\rightarrow\infty$ and $z\rightarrow 0^{+}$).
@@
$f_{n}^{\prime}(z)$ $={\mathrm{tr}}\bigg{[}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}{\bm{\Sigma}}\bigg{(}\frac{{\bm{\Sigma}}^{\frac{1}{2}}\mathbf{Z}^{\top}\mathbf{Z}{\bm{\Sigma}}^{\frac{1}{2}}}{n}+\mathbf{Q}+z{\bm{\Sigma}}\bigg{)}^{-1}\bigg{]}\mathbf{A}.$
By Assumption 2.5, when the constant $c_{2}\not=0$, it is easy to bound $f^{\prime}_{n}(z)$ by $\frac{\|{\bm{\Sigma}}^{-1/2}\mathbf{A}{\bm{\Sigma}}^{-1/2}\|_{{\mathrm{tr}}}}{c^{\prime 2}c_{2}^{2}}$ with some constant $c^{\prime},c_{2}>0$. We then consider the case $c_{2}=0$ and $c^{\prime}\mathbf{I}\leq\mathbf{Q}\leq C^{\prime}\mathbf{I}$. If we can prove that for any fixed vector $\mathbf{u}$ with $\|\mathbf{u}\|=1$,
@@
almost surely for large enough $n$. Here, $\lambda_{\mathbf{A},i}$ is the eigenvalue of matrix $\mathbf{A}$ and ${\bm{\xi}}_{\mathbf{A},i}$ is the corresponding eigenvector. Hence, the main target now is to deduce the uniform bound of (C.6). We separate the proof of the uniform bound of (C.6) into the case $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$ and $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$.
When $z\geq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$, it can be seen that (C.6) is smaller than the value when we let $z=\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$, hence it remained for us to complete the proof when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$.
When $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$, by Assumption 2.6, the number of eigenvalues tending to infinity remains fixed, we can separate ${\bm{\Sigma}}$ by
@@
Therefore, when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$, we have almost surely
$\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}\leq\|\mathbf{Q}\|_{{\mathop{\mathrm{op}}}}+(z\cdot\sum_{j=1}^{K}\sqrt{3c\lambda_{j}})=O(1).$
for $n$ large enough. Hence when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$,
@@
almost surely for all $n$ large enough. Here, the first inequality utilizes ${\widehat{\lambda}}_{K}\geq 2\lambda_{K}/3$ almost surely for large enough $n$. We prove that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$ for any $j\in[1:K]$,
$\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\cdot\sqrt{\lambda_{j}}=\sqrt{\lambda_{j}}/\lambda_{\min}(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{B}^{\top})\leq 2\sqrt{\lambda_{j}}/\lambda_{K}\leq 2C.$ (C.11)
almost surely for all $n$ large enough. Note that $\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}=O(1)$ as long as $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$ for $n$ large enough, we conclude that
$\|{\bm{\Delta}}_{bs}\|_{{\mathop{\mathrm{op}}}}\cdot\sqrt{\lambda_{j}}\leq\sqrt{\lambda_{j}}\|{\bm{\Delta}}_{ss}\|_{{\mathop{\mathrm{op}}}}\|\mathbf{B}\|_{{\mathop{\mathrm{op}}}}\leq 3C$ (C.12)
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$ for some large enough absolute $C>0$. As for $\|{\bm{\Delta}}_{bb}\|_{{\mathop{\mathrm{op}}}}$, we have
@@
almost surely for all $n$ large enough when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$ Combined (C.9) with (C.10)-(C.13), we conclude that when $z\leq\sqrt{\frac{\lambda_{\min}(\mathbf{Q})}{18c(\sum_{j=1}^{K}\sqrt{\lambda_{j}})^{2}}}$,